OpenStudy (anonymous):

What is the direct linear variation equation for the relationship? y varies directly with x and y = 4 when x = 12. A. y = x + 8 B. y = 3x C. y = x – 8 D. 01_15_UT_07a.gif

OpenStudy (anonymous):

@iGreen @KendrickLamar2014

OpenStudy (anonymous):

@Joel_the_boss @iambatman

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@77777jeannie77777 @BlossomCake

OpenStudy (anonymous):

@bohotness @confluxepic

OpenStudy (anonymous):

@Hopkiand002

OpenStudy (campbell_st):

a direct variation equation is in the form y = kx k is called the variation constant so substitute x = 12 and y = 4 to find the value of the constant and then you'll have your equation

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

which would be k? @campbell_st

OpenStudy (campbell_st):

you can caclulate k... by substituting the given values of x = 12 and y = 4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

solve \[4 = k \times 12\]

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

@campbell_st is it b?

OpenStudy (campbell_st):

nope... and I can't see D direct variation means the y intercept is zero... so that eliminates A and C

OpenStudy (anonymous):

d is y=1/3x

OpenStudy (campbell_st):

so using the equation I posted k = 4/12 simplify the fraction.